OpenStudy (anonymous):

Provide the solution to the system - x + 4y = –16 3x + 2y = 12

OpenStudy (mertsj):

(8,-6)

OpenStudy (slaaibak):

Write the first equation in terms of y: x = - 16 - 4y Now substitute that x value in the second equation's x

OpenStudy (anonymous):

x + 4y = –16 3x + 2y = 12 it's not neg. lol that's why i wasnt getting the right answer. my bad sorry :///

OpenStudy (mathmate):

To Alice: You have submitted more than 30 problems of the same type (linear equations). The idea of the exercises is to give you practice. I feel bad in that we are the ones benefiting from the exercises. It would be beneficial to your learning if you attempt some problems by yourself. What do you think? MathMate.